### v1

# Learning Penalty for Optimal Partitioning via Automatic Feature Extraction

Abstract
Changepoint detection identifies significant shifts in data sequences, making it important in areas like finance, genetics, and healthcare. The Optimal Partitioning algorithms efficiently detect these changes, using a penalty parameter to limit the changepoints number. Determining the appropriate value for this penalty can be challenging. Traditionally, this process involved manually extracting statistical features, such as sequence length or variance to make the prediction. This study proposes a novel approach that uses recurrent neural networks to learn this penalty directly from raw sequences by automatically extracting features. Experiments conducted on 20 benchmark genomic datasets show that this novel method surpasses traditional methods in partitioning accuracy in most cases.
keywords:
changepoint detection, optimal partitioning, penalty learning, supervised machine learning, recurrent networks
[…]
[xxx]organization=Northern Arizona University, city=Flagstaff, state=Arizona, country=United States
[…]
[yyy]organization=Université de Sherbrooke, city=Quebec, country=Canada
1 Introduction
Changepoint detection (also known as partitioning or segmentation) is crucial for identifying abrupt changes in data in various domains, including finance (Lattanzi and Leonelli, 2021), healthcare (Muggeo and Adelfio, 2010), network security (Tartakovsky et al., 2013), and environmental science (Reeves et al., 2007). This importance has driven the development of numerous methods over the past several decades. The literature review, gap in existing methods, and study contributions are based on the diagram in Figure 1.
Figure 1: Literature review of the changepoint detection process in this study: (i) Selecting a detection method, (ii) Choosing a regularization method to control the changepoints number, (iii) Selecting a model type to predict the penalty, (iv) Set up the penalty prediction supervision, and (v) Choosing the model architecture
(i) (ii) Related Changepoint Detection Methods.
Changepoint detection methods can be divided into two categories: Deep Learning (DL) and Dynamic Programming (DP).
DL-based methods, which have evolved from earlier method Cumulative Sum (CUSUM) (Hinkley, 1971; James et al., 1987), including recent works (Li et al., 2024; Ermshaus et al., 2023) employ multilayer perceptrons (MLP) or k-nearest neighbors (KNN) as classification sliding windows to identify changepoints. However, these methods rely on local neighbors rather than the entire sequence and are overly complex for univariate data.
In contrast, DP is well-suited for this context. These algorithms rely on a single hyperparameter to limit the changepoints number, which can be either: (a) a fixed changepoints number or (b) a penalty parameter to penalize the changepoint presence. Fixed changepoints number allow efficient DP algorithms (Auger and Lawrence, 1989; Bai and Perron, 2003; Killick et al., 2012; Rigaill, 2015) to locate them, but the changepoints number is rarely predetermined. More commonly used, penalty-based DP algorithms, such as Optimal Partitioning (OPART) (Jackson et al., 2005) and its variants—Pruned Exact Linear Time (PELT) (Killick et al., 2012), Functional Pruning Optimal Partitioning (FPOP) (Maidstone et al., 2016), and Labeled Optimal Partitioning (LOPART) (Hocking and Srivastava, 2023)—are widely regarded as the most effective. PELT and FPOP prune candidate changepoints efficiently, producing identical partitions as OPART, while LOPART extends OPART by incorporating predefined labels (the expected number of changepoints within specific location ranges) or defaulting to OPART operation when labels are absent.
In OPART family algorithms, the penalty parameter plays a crucial role. A higher penalty imposes a stronger penalty on changepoint presence, leading to fewer detected changepoints, see Figure 2. While OPART use a fixed penalty, this study aims to predict this penalty value to enhance OPART accuracy. From here on, the OPART penalty will be referred to as $\lambda$. To gain a detailed understanding of how $\lambda$ works, please refer to the subsection “OPART Optimization Problem” in the appendix.
Figure 2: Example demonstrating how different $\lambda$ values affect the OPART detection of changepoints. The labeled sequence contains four labels: three positive (one changepoint regions) and one negative (no changepoints region). Various $\lambda$ are tested in the OPART algorithm. Two types of label errors are considered: false positives (fp), where extra changepoints are detected in positive labels or any in negative labels, and false negatives (fn), where no changepoint is detected in a positive label.
(iii) (iv) Existing $\lambda$ prediction models limitation.
Various models exist for predicting $\lambda$, including unsupervised (Schwarz, 1978; Akaike, 1974; Lavielle, 2005) and supervised such as linear (Rigaill et al., 2013) and trees (Drouin et al., 2017; Barnwal et al., 2022). But all of these models rely on a set of manually extracted statistical sequence features, yet technically the number of possible features is infinite, this process can result in the inclusion of many irrelevant features, and the potential omission of useful ones.
(iv) (v) Contribution.
This study introduces a novel approach utilizing a recurrent neural network to automatically extract relevant features from sequences, which are then used to predict $\lambda$. Experimental results on benchmark datasets show that this approach surpasses previous methods in terms of OPART accuracy in the most cases.
Reproducibility.
To ensure transparency and fairness, the code used to generate the results of this study is accessible at: https://github.com/lamtung16/ML_Changepoint_Detection_epigenomic_rnn.
2 Problem Setting and Previous Work
2.1 Problem Setting
Study supervision.
Each sequence is assigned labels, each indicating the expected changepoints count in a specific region, see the example in Figure 2 of a sequence having four labels. DL-based methods are not applicable here because they require supervision with exact changepoint locations.
The prediction models use interval regression.
The objective of this study is to predict the optimal $\lambda$ for each sequence to detect its changepoints. For each labeled sequence, the optimal $\lambda$ is not a single value, but rather exists within an interval $[\lambda_{l},\lambda_{u}]$. For example, in Figure 2, $\lambda=10$ is the optimal value, as it generates changepoints that align with expert labels; however, there exist $a,b\geq 0$ such that values of $\lambda$ within $[\lambda_{l},\lambda_{u}]=[10-a,10+b]$ could produce the same result. The algorithm for generating the optimal interval for each labeled sequence is detailed by Rigaill et al. (2013). This problem, where the prediction needs to fall within an interval, is called interval regression. In interval regression problem, there are 4 types of intervals: uncensored $(-\infty<\lambda_{l}=\lambda_{u}<\infty)$, interval-censored $(-\infty<\lambda_{l}<\lambda_{u}<\infty)$, left-censored $(-\infty=\lambda_{l}<\lambda_{u}<\infty)$ and right-censored $(-\infty<\lambda_{l}<\lambda_{u}=\infty)$.
2.2 Previous Work
This subsection focuses on the process of training supervised $\lambda$ prediction models. Before delving into the details, some noteworthy models are briefly highlighted, which, although not directly applicable to this study, are still relevant. Additionally, some unsupervised models, which have served as inspiration for supervised models, are mentioned.
Noteworthy.
Adaptive Linear Penalty INference (ALPIN) (Truong et al., 2017) is a notable method for a special case of this supervision, where every single point in the sequence is a label, equivalent to label regions having length 1. However, since this study uses a more relaxed supervision with varying label region lengths, ALPIN is not applicable.
The Accelerated Failure Time (AFT) model (Wei, 1992) and its variants (Cai et al., 2009; Huang et al., 2006; Quinlan, 1986; Pölsterl et al., 2016), which are designed for censored outcomes, can be considered. Supervised changepoint detection setups involve interval-censored targets, whereas traditional AFT models are restricted to uncensored or right-censored intervals, making them not directly applicable. An important distinction is that in AFT models, the left-censored interval is defined as $(0=\lambda_{l}<\lambda_{u}<\infty)$, which is different from this study’s definition.
Unsupervised.
The unsupervised models predicts $\lambda$ based on some sequence statistical features. Let $N$ denote the length and $\sigma$ the standard deviation of the sequence, with examples such as Baysian Information Criterion (BIC) (Schwarz, 1978) predicts $\lambda=\log N$ or $\lambda=\sigma^{2}\log N$. Akaike’s Information Criterion (AIC) (Akaike, 1974) predicts $\lambda=2p$, where $p$ represents a sequence feature (e.g., standard deviation, variance). Lavielle (2005), on the other hand, does not provide a closed-form formula but considers $\lambda$ to be dependent on both $N$ and $\sigma$.
The rest of this section will detail the training process of three supervised machine learning models: Maximum Margin Interval Regression (Rigaill et al., 2013), a linear model; Maximal Margin Interval Tree (MMIT) (Drouin et al., 2017); and AFT in XGBoost (Barnwal et al., 2022), which utilizes a tree-based framework (although AFT in XGBoost is an AFT model, it is applicable due to its ability to handle all types of non-negative intervals). These models divided the training process into 2 independent steps:
  • 1.
    Step 1: Extract features vector $\mathbf{x}$ from the sequence.
  • 2.
    Step 2: Train a model $g(\cdot)$ takes into $\mathbf{x}$ to predict $\lambda$, $\lambda=g(\mathbf{x})$
2.2.1 Step 1: Feature Extraction
The manual feature extraction process, detailed in Rigaill et al. (2013), involves the following steps:
  • 1.
    Starting from the sequence $\mathbf{d}=[d_{1},d_{2},\dots,d_{N}]$ with the mean $\bar{d}=\frac{1}{N}\sum_{i=1}^{N}d_{i}$, 2 additional vectors (residuals and differences) are generated:
    $\mathbf{d}_{\text{res}}$ $=[d_{1}-\bar{d},\dots,d_{N}-\bar{d}]\in\mathbb{R}^{N}$
    $\mathbf{d}_{\text{diff}}$ $=[d_{2}-d_{1},\dots,d_{N}-d_{N-1}]\in\mathbb{R}^{N-1}$
  • 2.
    Apply 3 transformations (identity, absolute, square) to each of 3 vectors, resulting in 9 transformed vectors.
  • 3.
    Compute 8 statistics (sum, mean, standard deviation, and quantiles at 0%, 25%, 50%, 75%, 100%) from each transformed vector, yielding a feature vector of length 72.
  • 4.
    Include the sequence length as an additional feature, forming a vector of length 73.
  • 5.
    Apply 5 transformations (identity, square root, log, log-log, square) to each feature, resulting in up to 365 features.
Undefined values, such as those caused by invalid operations (e.g., log of a negative number), is ignored.
2.2.2 Step 2: Model Training
After obtaining the feature vector $\mathbf{x}$ (up to 365 features), a model is selected and trained to predict $\lambda$.
AFT in XGBoost.
The model is formulated as:
$\lambda=\mathcal{\bar{T}}(\mathbf{x})+\epsilon$
where $\mathcal{\bar{T}}(\mathbf{x})$ is the output from the decision tree ensemble based on the feature vector $\mathbf{x}$, and $\epsilon$ is a random error with a specified distribution (e.g., normal, log-normal, Weibull). To train the model, the error distribution is selected, and the AFT likelihood is constructed. Maximum likelihood estimation is then used for training. Optimal hyperparameters are chosen through cross-validation on the train set.
Figure 3: Example about the loss value of two loss functions is determined by comparing the prediction to the target. In the top plot, the Squared Error yields a value of 0 when the prediction $\hat{y}$ reaches the target point $y$. The bottom plot illustrates the Squared Hinge Error, where the loss value reaches 0 when the prediction $\hat{y}$ falls within the interval $[y_{l}+\epsilon,y_{u}-\epsilon]$
Linear.
This model, known as Maximum Margin Interval Regression, transforms the problem to predicting $\log\lambda$ instead of $\lambda$, as $\log\lambda$ can take any real value, unlike the non-negative $\lambda$. The model expresses the output as a linear combination of inputs:
$\log\lambda=\mathbf{x}\cdot\beta+\beta_{0}$
where $\beta$ is the parameter vector, $\mathbf{x}\cdot\beta$ is the dot product of $\mathbf{x}$ and $\beta$, and $\beta_{0}$ is the bias term. Its simplified version used in Hocking and Srivastava (2023) predicts $\log\lambda=\beta_{1}\times\log\log N+\beta_{0}$, where $N$ is the sequence length.
This model is trained by minimizing a specific loss function. Squared error (SE), commonly used in point estimate regression, is not suitable for this problem since each target is not a single point. Instead, the Hinge Error loss function, defined by Rigaill et al. (2013), generated from SE, is used. Unlike SE, which achieves 0 loss at a single point (when the prediction is identical with the target point), Hinge Error achieves 0 loss within an interval (when the prediction falls within the target interval), as shown in Figure 3. Using the ReLU function, the Hinge Error between prediction $\hat{y}$ and target interval $[y_{l},y_{u}]$ is expressed as:
$l(\hat{y},[y_{l},y_{u}])=\Big{(}\text{ReLU}(y_{l}-\hat{y}+\epsilon)\Big{)}^{p}+\Big{(}\text{ReLU}(\hat{y}-y_{u}+\epsilon)\Big{)}^{p}$ (1)
Here, the margin length $\epsilon$ is fixed ($\epsilon=1$ by default) and the loss type $p=2$ (Squared Hinge Error) are chosen.
Tree.
MMIT introduced by Drouin et al. (2017) – similar to regression trees by Breiman et al. (1984) – is applicable for $\lambda$ prediction as an interval regression model. Unlike Breiman et al. (1984), which minimizes Squared Error, MMIT minimizes the Hinge Error 1 within tree nodes to construct the prediction tree model:
$\log\lambda=\mathcal{T}(\mathbf{x})$
Here, $\mathcal{T}$ is the tree model. The optimal tree architecture (maximum depth, minimum sample split), loss type which is $p$ (either 1 or 2) in Hinge Error 1, and margin which is $\epsilon$ in Hinge Error 1, is selected through cross-validation.
3 Novelty
In previous work, the prediction of $\lambda$ involves two independent steps. In step 1, which manually extracts sequence features, redundancy may be introduced, and important hidden features may be missed. To address this, we propose a model that combines both steps into one, utilizing a recurrent network to automatically extract relevant features from the raw sequence to predict $\lambda$, see Figure 4.
Figure 4: Diagram of the methods - Instead of performing feature extraction and the learning model separately, the proposed method uses a recurrent network to combine both into a single operation.
Feature Extraction via Recurrent Networks.
In a Recurrent Network with input size $k\geq 1$ and hidden size $m\geq 1$, processing a sequence $\mathbf{d}=[d_{1},d_{2},\dots,d_{N}]\in\mathbb{R}^{N}$, the hidden state at time $t$, called $h_{t}$, is updated as:
$h_{t}=f(d_{[t:t+k-1]};h_{t-1})\in\mathbb{R}^{m}$
where $d_{[t:t+k-1]}=[d_{t},d_{t+1},...,d_{t+k-1}]\in\mathbb{R}^{k}$. Time stamp $t$ ranges from $k$ to $N$. The initial hidden state $h_{k-1}$ can be assigned manually. The final hidden state $h_{N}\in\mathbb{R}^{m}$ captures information from the entire sequence. Thus, we use $h_{N}$ as the $m$ extracted features from the raw sequence.
Predicting $\lambda$ from extracted features.
The prediction is formed by the linear combination of $m$ extracted features:
$\log\lambda=h_{N}\cdot\beta+\beta_{0}$
where $\beta\in\mathbb{R}^{m}$ and $\beta_{0}\in\mathbb{R}$ are the parameters to be learned.
Figure 5: Example: visualization of features vs. targets from the dataset detailed. The upper plots show the relationship between length or variance and the target, as these features are considered key based on $\lambda$ prediction unsupervised methods BIC and AIC. The two lower plots illustrate the visualization of automatically extracted features vs. targets, using a GRU model with one hidden layer and a hidden size of 2, which corresponds to the number of extracted features. Since we consider that the predicted $\lambda$ has a linear relationship with the features, a good feature should exhibit an approximately linear relationship in the plot. In other words, a good feature should allow us to intuitively draw a straight line that above the lower targets (blue dots) and below the upper targets (red dots). From the figure, it is evident that Extracted Feature 1 is a better feature (compared to length or variance), as we can visually draw a clear line that separates the lower and upper targets.
Automatic extracted features vs. Manual extracted features.
The novel method generates useful features without the need for a large number, unlike previous models. Figure 5 provides an example by comparing the ability to predict $\lambda$ using length (or variance) - as these manual extracted features are considered key for $\lambda$ prediction in unsupervised methods - versus automatic extracted features from recurrent networks. It demonstrates that the automatic extracted features probably provide a more accurate prediction of $\lambda$ than either length or variance.
Feature Approximation Ability of Recurrent Networks.
Recurrent networks have the ability to approximate statistical sequence features. For example, if the core network is defined as $h_{t}=h_{t-1}+1$ when $h_{0}=0$ and the input size $k=1$, then $h_{N}=N$, which represents the sequence length, a feature used in BIC. The core component of these models is a deep neural network. For instance, the Vanilla Recurrent Network (Rumelhart et al., 1986) core is a MLP, which, according to the universal approximation theorem (Hornik et al., 1989), can approximate any continuous function, regardless of its complexity, provided there are enough neurons in the hidden layer. This is why, although these models may not directly produce exact statistical sequence features, their architecture enables them to approximate these features with a high degree of accuracy.
Recurrent networks to be implemented.
This study explores three recurrent networks: Vanilla Recurrent Network (RNN), Long Short-Term Memory (LSTM) (Hochreiter, 1997), Gated Recurrent Unit (GRU) (Chung et al., 2014). RNNs struggle with long-term dependencies due to the vanishing gradient problem. LSTMs address this by introducing memory cells and gating mechanisms that help retain information over longer sequences. GRUs, a simplified version of LSTMs, combine the forget and input gates into a single update gate, offering a more efficient alternative while still capturing long-term dependencies.
4 Experiments
In this section, the implementation of models is detailed to enhance reproducibility. Labeled sequences are first processed, with features manually extracted to replicate previous work, along with the corresponding target intervals. Baseline models and the proposed models are then implemented. For each test sequence, OPART is applied with the predicted $\lambda$ to detect changepoints. Finally, accuracy rates are calculated based on the predicted changepoints and the test set labels.
Table 1: List of datasets with its number of sequences (Size)
Except for the three datasets: systematic, detailed, and cancer, each ChIP-seq dataset is named using the following format: for example, ATAC_JV_adipose. The first part, “ATAC”, indicates the assay type. The second part, “JV”, represents the initials of the biologist who assigned the labels. The last part, “adipose”, refers to the sample set.
Dataset Seq. length Size Source
systematic   000066 - 5937 3418
detailed   000025 - 5937 3730 GitHub
cancer   000039 - 43628 826
ATAC_JV_adipose 105716 - 11.50M 465
CTCF_TDH_ENCODE 166431 - 11.09M 182
H3K27ac-H3K4me3_TDHAM_BP   000275 -0 4.33M 2008
H3K27ac_TDH_some 41053 -0 6.74M   0095
H3K27me3_RL_cancer   005730 -0 1.89M 171
H3K27me3_TDH_some 113523 -0 3.00M   0043
H3K36me3_AM_immune 184092 -0 7.14M 420
H3K36me3_TDH_ENCODE 68209 -0 2.74M   0078
H3K36me3_TDH_immune 434882 -0 4.35M   0037 UCI Repo
H3K36me3_TDH_other 49362 -0 7.15M   0040
H3K4me1_TDH_BP 148275 -0 9.73M 144
H3K4me3_PGP_immune 24460 -0 6.62M 297
H3K4me3_TDH_ENCODE 18198 -0 7.21M   0075
H3K4me3_TDH_immune 87557 -0 7.81M 378
H3K4me3_TDH_other 80674 -0 3.72M   0090
H3K4me3_XJ_immune 52498 -0 5.50M 270
H3K9me3_TDH_BP 369654 - 10.29M 120
Raw Sequence Datasets.
This study uses 20 datasets: two DNA copy number profiles from neuroblastoma tumors (Rigaill et al., 2013), known for detailed and systematic dataset; one dataset cancer from (Hocking and Srivastava, 2023); and 17 ChIP-seq datasets from (Hocking et al., 2016), see Table 1 for all benchmark datasets.
Train/Test Setup.
Each sequence is assigned a unique identifier (sequenceID), and datasets are split into folds based on these IDs. Fewer folds are used for smaller datasets to balance train and test data sizes, ensuring a sufficiently large test set for reliable evaluation. In each iteration, one fold serves as the test set, and the others form the train set.
Evaluation Metrics.
The evaluation metric used is the accuracy rate on the test set, using the label errors number and the total labels number.
4.1 Model Implementation
Previous models were implemented as baselines, along with additional models and the proposed models for comparison. Details of all models are provided in Table 2.
Table 2: List of employed models
Model Regularization Citation
linear L1 Rigaill et al. (2013)
MMIT
max depth
min split sample
loss type
loss margin
Drouin et al. (2017)
AFT_XGboost
learning rate
max depth
min child weight
reg alpha
reg lambda
Barnwal et al. (2022)
constant none Baseline
MLP
hidden layer number
hidden layer sizes
Additional
RNN
LSTM
GRU
hidden layer number
hidden state sizes
Proposed
4.1.1 Baseline Models
Linear.
R package penaltyLearning (Hocking, 2024) was used to implement the Max Margin Interval Regression model with L1 regularization. The regularization parameter L1 starts at 0.001, increasing by a factor of 1.2 until no features remain, with cross-validation determining the optimal L1 regularization value.
MMIT.
The MMIT model was implemented using the R package mmit (Drouin, 2017). Cross-validation is used to select optimal hyperparameters, including:
  • 1.
    max_depth: values of 0, 1, 5, 10, 20, and $\infty$
  • 2.
    min_sample: values of 0, 1, 2, 4, 8, 16, and 20.
  • 3.
    margin ($\epsilon$ in Hinge Error 1): values of 0, 1, and 2
  • 4.
    loss_type: values of hinge and square (equivalent to $p=1$ and $p=2$ in Hinge Error 1);
AFT in XGBoost.
The model is implemented using the Python package xgboost. Following the same setup described in Barnwal et al. (2022), cross-validation was performed to select hyperparameters, including:
  • 1.
    learning_rate: 0.001, 0.01, 0.1, and 1.0
  • 2.
    max_depth: 2, 3, 4, 5, 6, 7, 8, 9, and 10
  • 3.
    min_child_weight: 0.001, 0.1, 1.0, 10.0, and 100.0
  • 4.
    reg_alpha: 0.001, 0.01, 0.1, 1.0, 10.0, and 100.0
  • 5.
    reg_lambda: 0.001, 0.01, 0.1, 1.0, 10.0, and 100.0
  • 6.
    aft_loss_distribution_scale: 0.5, 0.8, 1.1, 1.4, 1.7, and 2.0
Constant (Featureless).
This model predicts a single value $\lambda^{*}$ from the train set using only target intervals. The objective is to minimize the total Hinge Error over $n$ instances with target intervals $[\lambda_{l}^{i},\lambda_{u}^{i}]_{i=1}^{n}$, defined as:
$\mathcal{L}(\lambda)=\sum_{i=1}^{n}l(\lambda,[\lambda_{l}^{i},\lambda_{u}^{i}]),\quad\lambda^{*}=\arg\min_{\lambda}\mathcal{L}(\lambda).$
4.1.2 Additional Model
MLP.
Using the same feature vector and loss function as the linear model, the MLP generalizes the linear model by incorporating ReLU activation in the hidden layers. Hyperparameters include the number of hidden layers (1, 2, or 3) and the size of each hidden layer (1, 2, 4, 8, 16, 32, 64, 128, 256, or 512). The best hyperparameters are chosen via cross-validation in the train set. Since the linear model uses L1 regularization and MMIT (or AFT in XGBoost) has a tree-based architecture, they can automatically select features. However, MLP does not fully mitigate irrelevant features on its own (parameters associated with each feature are likely non-zero). Feature selection for MLP in this interval regression setting could be explored in future work.
4.1.3 Proposed Models
Architecture.
There are three considered models: RNN, LSTM, and GRU. The hyperparameters include: the number of hidden layers (1 or 2) and the hidden state size (number of extracted features) (2, 4, 8 or 16).
Preprocessing.
Based on the sequence lengths of 20 datasets, we handle them as follows:
  • 1.
    For the 3 datasets—detailed, systematic, cancer—the sequence lengths range from 39 to 43628. These sequences are directly input into the $\lambda$ prediction models without further modification.
  • 2.
    For the 17 ChIP-seq datasets, the sequence lengths can be over 11 million. Training on these raw sequences can be computationally expensive and time-consuming. So pooling operations are performed, where a single representative value is chosen for each non-overlapping sliding window. The window size options are 100, 1000, or 2000, and the representative value is either the window mean or the median. After shortening the sequences, a $\log(z+1)$ transformation is applied to all value $z$ in all shortened sequences to mitigate skewed distributions. Following this, all sequence values are normalized (mean 0, variance 1) for each dataset.
Figure 6: The accuracies for each fold are presented for datasets without any pooling preprocessing. The cancer dataset has 2 folds, while the other two datasets each have 6 folds. The Constant and RNN models have been excluded due to their small accuracies to better highlight the differences and improve visualization. Overall, the GRU model achieves the highest accuracy.
Figure 7: The fold accuracies on some ChIP-seq datasets are presented, with the datasets ordered from largest to smallest. The number of colored diamonds in any row represents the number of folds for each dataset. As observed, for larger datasets, more complex models tend to show higher accuracy. For the last three smaller datasets, recurrent networks do not always outperform the previous methods.
4.2 Results
Figure 6 presents the accuracy rates for three datasets using the original, unpooled sequences, while Figure 7 shows the results for various ChIP-seq datasets (pooled sequences). Each dataset is evaluated using seven different methods, producing multiple accuracy values depending on the number of folds used for that dataset.
5 Discussion and Conclusion
Summary of Contributions.
Technically, these proposed models still perform two steps: (1) Feature extraction and (2) Predicting $\lambda$ based on the extracted features. However, instead of performing these steps separately, it combines them into a single operation, enhancing its robustness compared to previous models. Automatic extracting features helps generate highly useful ones without requiring many features, as shown in Figure 5. This study focuses exclusively on genomic datasets, where prior knowledge of relevant features enables effective manual feature extraction. However, the proposed models stand out because it does not require expert knowledge to perform. Its automatic feature extraction mechanism makes it versatile and applicable to any type of sequence dataset.
Comparison to Previous Models.
Recurrent networks, in general, outperform previous models across datasets, see Figure 6 and Figure 7. However, for smaller datasets (with fewer than 100 instances), recurrent networks and MLPs do not consistently outperform baseline models. On the other hand, when dealing with larger datasets, recurrent networks noticeably outperform previous models and show a slight edge over MLPs.
Comparison between Recurrent Networks.
When comparing RNN, LSTM, and GRU, as expected, RNN performs the worst due to its difficulty with long-term memory. In most cases, GRU outperforms the others.
Figure 8: Log of the loss validation with 68.27% confidence interval across different numbers of extracted hidden layers in the three proposed models shows that the number of extracted hidden features needs to be sufficient. As seen in the figure, when the number of features is 8 or 16, the loss value is lower compared to when the number of features is 2 or 4.
Discussion on the Number of Automatic Extracted Features.
As shown in the Figure 8, the number of automatic extracted features needs to be sufficiently large. It is observed that when 8 or 16 features are extracted, the validation loss value is slightly smaller compared to when only 2 or 4 features are extracted.
Figure 9: Log of the loss validation with 68.27% confidence interval across different pooling methods in the three proposed models reveals that the window size must be large enough to reduce training time and difficulty, yet small enough to retain crucial information. The figure shows that mean pooling outperforms median pooling. These results are based on 17 ChIP-seq datasets.
Impact of pooling preprocessing.
Figure 9 shows that shortening the sequence for training is effective. As the window size increases, more information is lost. However, when the window size is small, such as 100, the performance of the recurrent models is worse compared to larger window sizes like 1000 or 2000. This could be because the pooled sequence with a window size of 100 is still too long for the recurrent models to handle effectively. The pooling process is beneficial in two ways: it can improve the accuracy of recurrent networks and also reduce training time. We employed two pooling functions, mean and median, to represent a single value for each non-overlapping window. As shown in Figure 9, it is evident that using the mean value as the representation consistently yields smaller loss.
Study Limitations.
Although the Recurrent Network architecture is straightforward, as its core essentially a deep neural network, training these models can be time-consuming due to the inherently sequential nature of the process. For some train/test pairs, even after applying pooling to reduce the sequence length, training a model with a maximum of 2 layers, each containing no more than 16 neurons, and a maximum of 1000 iterations, took over approximately 10 days, despite utilizing the powerful NVIDIA A100 GPU. These models also require significant memory to store the hidden states at each time step. To avoid out-of-memory (OOM) errors, it is crucial to allocate sufficient memory, especially for long sequences.
Future Directions.
For these particular genome datasets, it is assumed that $\lambda$ exhibits an increasingly monotonic behavior with respect to both sequence length and variance, which are employed in unsupervised methods. One approach is to leverage monotonic network architectures, such as min-max networks (Sill, 1997), or more advanced architectures like Lattice Networks (You et al., 2017), with modifications to the loss function for each linear region.
We can consider CNNs (LeCun et al., 1998) because they are capable of handling sequences of varying lengths and generating a single output by utilizing convolution layers to extract features across the sequence. Global pooling layers (e.g., Global Max or Average Pooling) aggregate these features into a fixed-size representation. This vector is then passed through a MLP to produce the final output.
To build upon the initial concept of utilizing recurrent networks in this study, various types of models could be explored in the future:
  • 1.
    Using a more complex model: This study predicts $\lambda$ by using a linear combination of the extracted features. Alternatively, we could consider using a more complex model such as MLP that processes the extracted features instead of relying on a linear combination.
  • 2.
    Attention mechanisms (Bahdanau et al., 2014) dynamically focus on relevant input parts, assigning importance weights to subsequences. Unlike sequential RNNs, they effectively capture long-range dependencies and handle varying input lengths, making them ideal for this task.
  • 3.
    Bidirectional Recurrent Networks (Schuster and Paliwal, 1997): The models implemented in this study are unidirectional, leading to challenges with long-term memory, since the later parts of the sequence tend to be more important than the earlier ones. Using a bidirectional recurrent network architecture can aim to address this limitation, enabling the model to better capture information from the sequence.
Declaration of generative AI technologies in the writing process
The author(s) utilized OpenAI’s ChatGPT to help rephrase sentences for improved clarity during the preparation of this work. All content was subsequently reviewed and edited by the author(s), who take full responsibility for the final published material.
Funding sources
This research did not receive any specific grant from funding agencies in the public, commercial, or not-for-profit sectors.
Appendix A Optimal Partitioning Optimization Problem
Given a sequence $\mathbf{d}\in\mathbb{R}^{N}$ and penalty parameter $\lambda\geq 0$, find the partition vector $\mathbf{m}\in\mathbb{R}^{N}$ that minimizes the following cost function:
$C_{\lambda}(\textbf{d},\textbf{m})=\sum_{i=1}^{N}l(d_{i},m_{i})+\lambda\sum_{i=1}^{N-1}I[m_{i}\neq m_{i+1}]$
Here, $l(m_{i},d_{i})$ represents the squared error between the sequence value $d_{i}$ and the mean segment $m_{i}$, i.e., $l(d_{i},m_{i})=(m_{i}-d_{i})^{2}$. The indicator function $I[m_{i}\neq m_{i+1}]$ equals 1 if a changepoint occurs ($m_{i}\neq m_{i+1}$) and 0 otherwise. Therefore, $\sum_{i=1}^{N-1}I[m_{i}\neq m_{i+1}]$ denotes the total number of changepoints.
Appendix B Statistics on the accuracy of each dataset with respect to each method
Table 3: Accuracy means and standard deviations (by method)
Dataset constant linear mmit AFT_XGboost mlp rnn lstm gru
cancer 63.37 ± 14.93 79.55 ± 0.70 75.97 ± 2.90 71.70 ± 4.10 78.62 ± 0.04 75.42 ± 2.12 77.29 ± 0.52 79.58 ± 1.96
detailed 67.45 ± 6.42 93.60 ± 1.07 93.80 ± 1.47 93.91 ± 1.20 94.95 ± 0.49 86.26 ± 4.64 93.26 ± 0.91 95.32 ± 0.53
systematic 56.87 ± 9.87 96.81 ± 0.46 97.16 ± 0.72 97.37 ± 0.75 97.92 ± 0.80 72.87 ± 29.01 96.64 ± 0.32 98.19 ± 0.47
ATAC_JV_adipose 75.05 ± 0.19 80.75 ± 0.75 78.74 ± 1.38 79.78 ± 0.38 81.55 ± 0.28 80.46 ± 0.33 81.16 ± 0.17 81.91 ± 0.32
CTCF_TDH_ENCODE 75.47 ± 1.71 84.15 ± 0.40 87.65 ± 0.54 93.32 ± 0.84 93.96 ± 0.38 86.58 ± 0.23 89.97 ± 2.50 91.52 ± 7.80
H3K27ac-H3K4me3_TDHAM_BP 74.93 ± 3.60 78.00 ± 1.22 85.06 ± 5.06 91.95 ± 2.41 91.76 ± 1.77 85.94 ± 1.78 89.33 ± 2.44 93.65 ± 1.80
H3K27ac_TDH_some 58.69 ± 8.35 80.23 ± 0.60 76.51 ± 2.48 72.06 ± 1.30 79.21 ± 3.05 75.63 ± 4.03 $7$ 0.56 78.76 ± 1.49
H3K27me3_RL_cancer 54.86 ± 1.35 64.52 ± 5.54 65.90 ± 2.32 65.23 ± 0.78 67.58 ± 3.57 64.43 ± 3.32 65.49 ± 3.24 64.22 ± 4.08
H3K27me3_TDH_some 62.19 ± 2.89 85.48 ± 1.90 77.19 ± 6.68 85.40 ± 3.35 77.83 ± 5.26 85.43 ± 2.36 87.21 ± 2.26 86.66 ± 2.01
H3K36me3_AM_immune 91.85 ± 0.93 92.75 ± 1.33 88.29 ± 6.18 92.61 ± 0.15 91.72 ± 0.09 90.60 ± 1.06 93.44 ± 0.70 93.89 ± 0.92
H3K36me3_TDH_ENCODE 56.10 ± 7.42 77.51 ± 0.29 75.57 ± 5.05 75.61 ± 6.73 73.16 ± 3.25 72.43 ± 9.49 73.28 ± 8.28 73.00 ± 8.69
H3K36me3_TDH_immune 96.39 ± 1.96 87.78 ± 11.00 87.63 ± 0.32 90.13 ± 3.85 88.62 ± 8.01 95.67 ± 2.97 95.67 ± 2.97 95.67 ± 2.97
H3K36me3_TDH_other 77.00 ± 2.83 93.50 ± 2.12 79.00 ± 15.56 91.50 ± 0.71 70.50 ± 17.68 94.09 ± 2.28 94.09 ± 2.28 88.74 ± 5.28
H3K4me1_TDH_BP 78.63 ± 1.69 89.10 ± 2.37 87.02 ± 2.18 85.20 ± 1.68 89.43 ± 1.56 88.39 ± 1.63 88.79 ± 1.07 88.39 ± 1.63
H3K4me3_PGP_immune 59.73 ± 5.00 71.79 ± 3.04 67.76 ± 3.35 70.11 ± 1.27 72.39 ± 0.45 70.68 ± 2.95 72.43 ± 1.89 73.55 ± 0.18
H3K4me3_TDH_ENCODE 92.33 ± 1.83 94.78 ± 1.04 95.37 ± 1.74 97.05 ± 1.25 93.89 ± 2.79 94.60 ± 1.62 94.94 ± 1.14 94.94 ± 1.14
H3K4me3_TDH_immune 68.85 ± 1.39 71.79 ± 3.04 67.76 ± 3.35 70.11 ± 1.27 80.91 ± 3.16 80.61 ± 1.50 81.42 ± 1.50 81.98 ± 1.29
H3K4me3_TDH_other 85.15 ± 4.73 71.79 ± 3.04 67.76 ± 3.35 70.11 ± 1.27 91.83 ± 0.65 89.74 ± 1.76 90.16 ± 1.39 89.90 ± 1.99
H3K4me3_XJ_immune 67.13 ± 0.89 69.44 ± 4.29 71.89 ± 2.29 74.00 ± 3.64 71.98 ± 4.33 70.98 ± 2.36 72.29 ± 3.41 75.06 ± 2.40
H3K9me3_TDH_BP 48.94 ± 1.50 84.29 ± 1.75 85.94 ± 0.11 79.39 ± 6.81 84.19 ± 2.81 86.16 ± 1.43 85.04 ± 1.62 86.08 ± 2.48